# Power Transform Revisited: Numerically Stable, and Federated

### v1

@@ Abstract
Power transforms are popular parametric techniques for making data more Gaussian-like, and are widely used as preprocessing steps in statistical analysis and machine learning. However, we find that direct implementations of power transforms suffer from severe numerical instabilities, which can lead to incorrect results or even crashes. In this paper, we provide a comprehensive analysis of the sources of these instabilities and propose effective remedies. We further extend power transforms to the federated learning setting, addressing both numerical and distributional challenges that arise in this context. Experiments on real-world datasets demonstrate that our methods are both effective and robust, substantially improving stability compared to existing approaches.
@@ 1 Introduction
Power transforms are widely used to stabilize variance and reduce skewness, and are described in textbooks as foundational tools for data preprocessing. Among them, the Box-Cox (Box and Cox,, 1964) and Yeo-Johnson (Yeo and Johnson,, 2000) transforms are the most prominent, and are frequently employed in statistical analysis and machine learning pipelines (Ruppert,, 2001; Fink,, 2009). Despite their popularity, directly implementing these mathematical functions leads to serious numerical instability issues, producing erroneous outputs or even causing program crashes. This is not a theoretical concern: as we show in Table 1, simple inputs can easily provoke such failures.
The issue has been identified in prior work such as the MASS package (Venables and Ripley,, 2002), but only partial solutions have been proposed. Recently, this issue was raised by Marchand et al., (2022). Unfortunately, their analysis includes some incorrect claims and unsound solutions. As shown in Figure 3, their method fails to identify optimal parameters and can crash even on simple datasets. A very recent study by Barron, (2025) also mentions numerical issues, but their remedy relies on simple replacement of the numerical function, which remains insufficient to ensure stability. Consequently, this foundational question has been unanswered until now.
In this paper, we provide a comprehensive analysis of numerical instabilities in power transforms. Our solution include log-domain computation, careful reformulation of critical expressions, and constraints on extreme parameter values, all of which are essential to ensure robust numerical behavior. Guided by our theoretical analysis, we also construct adversarial datasets that systematically trigger overflows under different floating-point precisions. Such data can occur naturally in real-world settings (e.g., Figure 11 in the Appendix) or be deliberately introduced by adversarial clients in federated learning scenarios.
@@
Extensive experiments on real-world datasets validate the effectiveness of our approach. Compared to existing implementations, including those of Marchand et al., (2022), our methods consistently achieve superior stability in both centralized and federated settings.
@@
    We propose remedies addressing each source of instability, yielding numerically stable algorithms in both centralized and federated learning settings.
@@ 2.1 Power Transform
The Box-Cox (BC) transformation (Box and Cox,, 1964) is a widely used power transform requiring strictly positive inputs ($x>0$):
@@
The Yeo-Johnson (YJ) transformation (Yeo and Johnson,, 2000) extends BC to accommodate both positive and negative values:
@@
The parameter $\lambda$ controls the shape of the transformation; when $\lambda\neq 1$, the mapping is nonlinear and alters the distribution. Figure 1 shows the transformation curves for various $\lambda$.
@@
where $\sigma^{2}_{\psi}=\mathrm{Var}[\psi(\lambda,x)]$. Both NLL functions are strictly convex (Kouider and Chen,, 1995; Marchand et al.,, 2022), guaranteeing a unique optimum. SciPy (Virtanen et al.,, 2020) estimates $\lambda^{*}$ using Brent’s method (Brent,, 2013), a derivative-free optimizer with superlinear convergence.
@@ 2.2 Numerically Stable Variance Computation
is well known to be numerically unstable (Ling,, 1974; Chan et al.,, 1983).
A numerically stable one-pass approach (Chan et al.,, 1982) maintains $(n,\bar{x},s)$ triples, where $s=\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}$. Merging two partial aggregates $(n^{(A)},\bar{x}^{(A)},s^{(A)})$ and $(n^{(B)},\bar{x}^{(B)},s^{(B)})$ is done as:
@@ 2.3 Federated Learning
Federated learning (Kairouz et al.,, 2021) is a distributed paradigm where multiple clients collaboratively train a model without sharing raw data, thereby preserving privacy. Each client computes local updates from its own data and transmits them to a central server, which aggregates the information to update the global model. Communication efficiency is a key concern, as excessive exchanges can be costly; methods such as FedAvg (McMahan et al.,, 2017) are commonly used to reduce communication overhead.
@@ 3 Understanding Numerical Instabilities
Numerical instabilities in power transforms arise in the computation of the NLL functions defined in Equations (3) and (4). These instabilities can disrupt optimization, producing suboptimal or even incorrect results. In practice, this means that the output of the method becomes unusable, causing bugs or failures in downstream tasks. Moreover, these issues cannot be resolved by simple quick fixes, as the instability is fundamental to the underlying computation and requires careful redesign for robust operation. As illustrated in Figure 3, the Exponential Search method (Marchand et al.,, 2022) [^1] There is a mismatch between the sign formula (Equation 3) and the ExpUpdate method (Algorithm 2) in (Marchand et al.,, 2022). The sign formula is $\partial\text{NLL}_{\text{YJ}}/\partial\lambda$ (not $\partial\ln\mathcal{L}_{\text{YJ}}/\partial\lambda$), so $\Delta=1$ in ExpUpdate should instead be $\Delta=-1$. Alternatively, one could add a negative sign in Equation 3 and keep ExpUpdate unchanged. fails even on simple datasets. The true optimum $\lambda^{*}$ can be quite large in magnitude, causing numerical overflow when evaluating the power functions, e.g., $0.1^{-361}$ or $10^{358}$. As a result, ExpSearch will crash and return values that diverge significantly from the true optimum.
@@
To better understand these issues, we first summarize key properties of the BC transformation $\psi(\lambda,x)$ in Theorem 3.1; the proof is intricate and deferred to Appendix A. Analogous results for YJ can be found in Yeo, (1997); Yeo and Johnson, (2000).
@@
    If $\psi^{(k)}=\partial^{k}\psi(\lambda,x)/\partial\lambda^{k}$ then, for $k\geq 1$.
@@
Avoiding Zero Points. By Theorem 3.1.1, $\psi(\lambda,1)=0$ for all $\lambda$. Thus, $x=1$ should be avoided. Choosing all $x>1$ ensures $\psi(\lambda,x)>0$ and may lead to positive overflow; similarly, choosing all $x<1$ can yield negative overflow. Datasets containing both $x>1$ and $x<1$ may be less likely to trigger overflow, and constructing simple adversarial examples in this regime is more challenging.
@@
Transf. Overflow Adversarial data $\lambda^{*}$ Extreme Value Max Value (Precision)
BC Negative [0.1, 0.1, 0.1, 0.101] -361.15 -3.87e+358 1.80e+308 (Double)
Positive [10, 10, 10, 9.9] 357.55 9.96e+354
YJ Negative [-10, -10, -10, -9.9] -391.49 -1.51e+407
Positive [10, 10, 10, 9.9] 393.49 1.51e+407
BC Negative [0.1, 0.1, 0.1, 0.10001] -35936.9 -2.30e+35932 1.19e+4932 (Quadruple)
Positive [10, 10, 10, 9.999] 35933.3 5.85e+35928
YJ Negative [-10, -10, -10, -9.999] -39524.8 -2.29e+41158
Positive [10, 10, 10, 9.999] 39526.8 2.292e+41158
BC Negative [0.1, 0.1, 0.1, 0.100001] -359353.0 -2.74e+359347 1.61e+78913 (Octuple)
Positive [10, 10, 10, 9.9999] 359349.0 6.99e+359343
YJ Negative [-10, -10, -10, -9.9999] -395283.0 -6.47e+411640
Positive [10, 10, 10, 9.9999] 395285.0 6.47e+411640
@@ 4 Numerically Stable Power Transform
This section presents a numerically stable approach for power transforms, addressing different sources of instability and then introduces remedies.
Modern computers use finite-precision floating-point arithmetic (e.g., IEEE 754 standard (IEEE,, 2019)), which limits the range of representable values. As shown in Table 1, power transforms can generate extreme values that exceed these limits, resulting in numerical overflow. Therefore, direct computation of the NLL functions is prone to instability.
Beyond the representation issue, the choice of optimization method also plays a critical role. In particular, derivative-based methods, such as Exponential Search (Marchand et al.,, 2022), are prone to instability, whereas derivative-free methods (e.g., Brent’s method) can achieve stable optimization. We detail these observations below.
@@
$\frac{\partial\text{NLL}_{\text{BC}}}{\partial\lambda}=\frac{1}{\sigma^{2}_{\psi_{\text{BC}}}}\Biggl(\sum_{i=1}^{n}\psi_{\text{BC}}(\lambda,x_{i})\cdot\psi^{(1)}_{\text{BC}}(\lambda,x_{i})-\frac{1}{n}\sum_{i=1}^{n}\psi_{\text{BC}}(\lambda,x_{i})\cdot\sum_{i=1}^{n}\psi^{(1)}_{\text{BC}}(\lambda,x_{i})\Biggr)-\sum_{i=1}^{n}\ln x_{i}$ (9)
@@
Derivative-free methods are stable. By contrast, derivative-free methods rely only on evaluating the NLL function itself, which can be stabilized. In particular, the NLL function involves only $\ln\sigma^{2}_{\psi}$, which can be represented in floating-point format and efficiently computed using log-domain methods (Haberland,, 2023) (see Appendix B).
@@
Equation (11) removes the constant $1/\lambda$, which avoids catastrophic cancellation (Weckesser,, 2019). For example, with $\lambda<-14$ and $x>1$, we have $x^{\lambda}\to 0$, and the subtraction $x^{\lambda}-1$ leads to severe precision loss. Therefore, computing the variance after the transformation becomes very unstable. The same occurs for large $\lambda$ with $x<1$. This instability, illustrated in the left panel of Figure 4, shows large fluctuations in the NLL curve when keeping the constant term, which will disrupts optimization whenever the evaluated $\lambda$ lies in such regions.
@@
Here, $\psi^{-1}_{\text{BC}}$ is the inverse Box-Cox transform, expressed via the Lambert $W$ function (Corless et al.,, 1996):
@@
The same approach applies to the Yeo-Johnson transformation (see Appendix C).
@@ 5 Federated Power Transform
To reduce communication rounds, one-pass variance computation is preferred. However, this approach introduces severe numerical instabilities, which directly affect both NLL evaluation and the optimization of $\lambda$. To illustrate this, we compare two approaches: the naive one-pass method (Equation (5), also used in Marchand et al., (2022)) and our numerically stable aggregation procedure (Algorithm 1). As shown in Figure 5, the naive method produces large fluctuations in the NLL curve, disrupting optimization. In contrast, our stable approach yields smooth curves and enables reliable optimization.
@@
Algorithm 2 Federated NLL Comptation (Box-Cox)
@@ 6 Empirical Evaluation
Our experiments contain two parts: (1) evaluating the effect of power transforms on downstream tasks, and (2) testing the numerical stability of our methods. Code is available at https://github.com/xuefeng-xu/powertf. All experiments were performed on an Apple MacBook M3 (16GB RAM), completing within 1 hour.
@@ 6.1 Downstream Effectiveness
We first evaluate the impact of power transforms on downstream classification tasks using three datasets: Adult (Becker and Kohavi,, 1996), Bank (Moro et al.,, 2012), and Credit (Yeh,, 2009). Dataset statistics are shown in Table 2. Each dataset is split into 80% training and 20% testing. Since features may contain negative values, we apply the Yeo-Johnson transformation to each feature, estimating $\lambda^{*}$ from the training set. We compare against two baselines: (1) Standardization (STD: zero mean and unit variance), and (2) Raw data without any transformation.
Table 2: Dataset statistics.
Datasets # Row # Col
Adult 33K 14
Bank 45K 16
Credit 30K 24
We then train three classifiers on the transformed features: Linear Discriminant Analysis (LDA), Quadratic Discriminant Analysis (QDA), and Logistic Regression (LR). Since LDA and QDA assume normally distributed inputs, power transforms are expected to improve performance. Figures 6 shows ROC curves and AUC scores on test set (additional plots are in Figure 9 in the Appendix). Power transforms consistently outperform the baselines, although the improvements are modest. A recent study by Eftekhari and Papyan, (2025) applied power transforms to hidden layers of deep neural networks, showing that improved Gaussianity leads to higher classification accuracy. This further confirms the effectiveness of power transforms.
@@
We next study the effect of varying $\lambda$ on AUC scores. This tests whether the estimated $\lambda^{*}$ is indeed optimal for downstream tasks. Since each dataset has multiple features, we select the feature with the highest mutual information with the label and ignore other features, apply the Yeo-Johnson transform with varying $\lambda$. Figures 7 shows AUC as a function of $\lambda$ (additional results are in Figure 10 in the Appendix).

### v2

@@ Abstract
Power transforms are popular parametric methods for making data more Gaussian-like, and are widely used as preprocessing steps in statistical analysis and machine learning. However, we find that direct implementations of power transforms suffer from severe numerical instabilities, which can lead to incorrect results or even crashes. In this paper, we provide a comprehensive analysis of the sources of these instabilities and propose effective remedies. We further extend power transforms to the federated learning setting, addressing both numerical and distributional challenges that arise in this context. Experiments on real-world datasets demonstrate that our methods are both effective and robust, substantially improving stability compared to existing approaches.
@@ 1 Introduction
Power transforms are widely used to stabilize variance and reduce skewness, and are described in textbooks as foundational tools for data preprocessing. Among them, the Box-Cox (Box and Cox, 1964) and Yeo-Johnson (Yeo and Johnson, 2000) transforms are the most prominent, and are frequently employed in statistical analysis and machine learning pipelines (Ruppert, 2001; Fink, 2009). Despite their popularity, directly implementing these mathematical functions leads to serious numerical instabilities, producing erroneous outputs or even causing program crashes. This is not a theoretical concern: as we show in Table 1, simple inputs can easily provoke such failures.
The issue has been identified in prior work such as the MASS package (Venables and Ripley, 2002), but only partial solutions have been proposed. Recently, this issue was raised by Marchand et al. (2022). Unfortunately, their analysis includes some incorrect claims and unsound solutions. As shown in Figure 4, their method fails to identify optimal parameters and can crash even on simple datasets. A very recent study by Barron (2025) also mentions numerical issues, but their remedy relies on simple replacement of the numerical function, which remains insufficient to ensure stability. Consequently, this foundational question has been unanswered until now.
In this paper, we provide a comprehensive analysis of numerical instabilities in power transforms. Our solution includes log-domain computation, careful reformulation of critical expressions, and constraints on extreme parameter values, all of which are essential to ensure robust numerical behavior. Guided by our theoretical analysis, we also construct adversarial datasets that systematically trigger overflows under different floating-point precisions. Such data can occur naturally in real-world settings (e.g., Figure 13 in the Appendix) or be deliberately introduced by adversarial clients in federated learning scenarios.
@@
Extensive experiments on real-world datasets validate the effectiveness of our approach. Compared to existing implementations, including those of Marchand et al. (2022), our methods consistently achieve superior stability in both centralized and federated settings.
@@
    We propose remedies addressing each source of instability, yielding numerically stable algorithms in both centralized [^1] Our implementation has been merged into SciPy. and federated learning settings.
@@ 2.1 Power Transform
The Box-Cox (BC) transformation (Box and Cox, 1964) is a widely used power transform requiring strictly positive inputs ($x>0$):
@@
The Yeo-Johnson (YJ) transformation (Yeo and Johnson, 2000) extends BC to accommodate both positive and negative values:
@@
The parameter $\lambda$ controls the shape of the transformation; when $\lambda\neq 1$, the mapping is nonlinear and alters the distribution. Figure 1 shows the transformation curves for various $\lambda$, while Figure 2 illustrates Box-Cox’s effect on left- and right-skewed data.
@@
Figure 2: Box-Cox transformation on left- (top) and right-skewed (bottom) data.
@@
where $\sigma^{2}_{\psi}=\mathrm{Var}[\psi(\lambda,x)]$. Both NLL functions are strictly convex (Kouider and Chen, 1995; Marchand et al., 2022), guaranteeing a unique optimum. SciPy (Virtanen et al., 2020) estimates $\lambda^{*}$ using Brent’s method (Brent, 2013), a derivative-free optimizer with superlinear convergence.
@@ 2.2 Numerically Stable Variance Computation
is well known to be numerically unstable (Ling, 1974; Chan et al., 1983).
A numerically stable one-pass approach (Chan et al., 1982) maintains $(n,\bar{x},s)$ triples, where $s=\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}$. Merging two partial aggregates $(n^{(A)},\bar{x}^{(A)},s^{(A)})$ and $(n^{(B)},\bar{x}^{(B)},s^{(B)})$ is done as:
@@ 2.3 Federated Learning
Federated learning (Kairouz et al., 2021) is a distributed paradigm where multiple clients collaboratively train a model without sharing raw data, thereby preserving privacy. Each client computes local updates from its own data and transmits them to a central server, which aggregates the information to update the global model. Communication efficiency is a key concern, as excessive exchanges can be costly; methods such as FedAvg (McMahan et al., 2017) are commonly used to reduce communication overhead.
@@ 3 Understanding Numerical Instabilities
Transf. Overflow Adversarial Data $\lambda^{*}$ NLL Extreme $\psi(\lambda^{*},x)$ Max Value
BC Negative [0.1, 0.1, 0.1, 0.109] -41.70 23.91 -1.20e+40 3.40e+38 (Single)
Positive [10, 10, 10, 9.2] 43.10 5.79 2.90e+41
YJ Negative [-10, -10, -10, -9.1] -40.10 5.32 -1.65e+42
Positive [10, 10, 10, 9.1] 42.10 5.32 1.65e+42
BC Negative [0.1, 0.1, 0.1, 0.101] -361.15 32.62 -3.87e+358 1.80e+308 (Double)
Positive [10, 10, 10, 9.9] 357.55 14.18 9.96e+354
YJ Negative [-10, -10, -10, -9.9] -391.49 14.18 -1.51e+407
Positive [10, 10, 10, 9.9] 393.49 14.18 1.51e+407
BC Negative [0.1, 0.1, 0.1, 0.10001] -35936.9 51.03 -2.30e+35932 1.19e+4932 (Quadruple)
Positive [10, 10, 10, 9.999] 35933.3 32.61 5.85e+35928
YJ Negative [-10, -10, -10, -9.999] -39524.8 32.61 -2.29e+41158
Positive [10, 10, 10, 9.999] 39526.8 32.61 2.292e+41158
BC Negative [0.1, 0.1, 0.1, 0.100001] -359353.0 60.24 -2.74e+359347 1.61e+78913 (Octuple)
Positive [10, 10, 10, 9.9999] 359349.0 41.82 6.99e+359343
YJ Negative [-10, -10, -10, -9.9999] -395283.0 41.82 -6.47e+411640
Positive [10, 10, 10, 9.9999] 395285.0 41.82 6.47e+411640
@@
Numerical instabilities in power transforms arise in the computation of the NLL functions defined in Equations (3) and (4). These instabilities can disrupt optimization, producing suboptimal or even incorrect results. In practice, this means that the output of the method becomes unusable, causing bugs or failures in downstream tasks. Moreover, these issues cannot be resolved by simple quick fixes, as the instability is fundamental to the underlying computation and requires careful redesign for robust operation. As illustrated in Figure 4, the Exponential Search method (Marchand et al., 2022) [^2] There is a mismatch between the sign formula (Equation 3) and the ExpUpdate method (Algorithm 2) in (Marchand et al., 2022). The sign formula is $\partial\text{NLL}_{\text{YJ}}/\partial\lambda$ (not $\partial\ln\mathcal{L}_{\text{YJ}}/\partial\lambda$); therefore, $\Delta=1$ in ExpUpdate should instead be $\Delta=-1$. Alternatively, one could add a negative sign in Equation 3 and keep ExpUpdate unchanged. fails even on simple datasets. The true optimum $\lambda^{*}$ can be quite large in magnitude, causing numerical overflow when evaluating the power functions, e.g., $0.1^{-361}$ or $10^{358}$. As a result, ExpSearch will crash and return values that diverge significantly from the true optimum.
@@
To better understand these issues, we first summarize key properties of the BC transformation $\psi(\lambda,x)$ in Theorem 3.1; the proof is intricate and deferred to Appendix A. Analogous results for YJ can be found in Yeo (1997); Yeo and Johnson (2000).
@@
    If $\psi^{(k)}=\partial^{k}\psi(\lambda,x)/\partial\lambda^{k}$ then, for $k\geq 1$:
@@
Staying on One Side of Zero Points. By Theorem 3.1.1, $\psi(\lambda,1)=0$ for all $\lambda$. Thus, $x=1$ should be avoided. Choosing all $x>1$ ensures $\psi(\lambda,x)>0$ and may lead to positive overflow; similarly, choosing all $x<1$ can yield negative overflow. Datasets containing both $x>1$ and $x<1$ may be less likely to trigger overflow, and constructing simple adversarial examples in this regime is more challenging.
@@ 4 Numerically Stable Power Transform
This section presents a numerically stable approach for power transforms, addressing different sources of instability and introducing corresponding remedies.
Modern computers use finite-precision floating-point arithmetic (e.g., IEEE 754 standard (IEEE, 2019)), which limits the range of representable values. As shown in Table 1, power transforms can generate extreme values that exceed these limits, resulting in numerical overflow. Therefore, direct computation of the NLL functions is prone to instability.
Beyond the representation issue, the choice of optimization method also plays a critical role. In particular, derivative-based methods, such as Exponential Search (Marchand et al., 2022), are prone to instability, whereas derivative-free methods (e.g., Brent’s method) can achieve stable optimization. We detail these observations below.
@@
$\frac{\partial\text{NLL}_{\text{BC}}}{\partial\lambda}=\frac{1}{\sigma^{2}_{\psi_{\text{BC}}}}\Biggl(\sum_{i=1}^{n}\psi_{\text{BC}}(\lambda,x_{i})\cdot\psi^{(1)}_{\text{BC}}(\lambda,x_{i})\\ -\frac{1}{n}\sum_{i=1}^{n}\psi_{\text{BC}}(\lambda,x_{i})\cdot\sum_{i=1}^{n}\psi^{(1)}_{\text{BC}}(\lambda,x_{i})\Biggr)-\sum_{i=1}^{n}\ln x_{i}$ (9)
@@
Derivative-free methods are stable. By contrast, derivative-free methods rely only on evaluating the NLL function itself, which can be stabilized. In particular, the NLL function involves only $\ln\sigma^{2}_{\psi}$, which can be represented in floating-point format and efficiently computed using log-domain methods (Haberland, 2023) (see Appendix B).
@@
Equation (11) removes the constant $1/\lambda$, which avoids catastrophic cancellation (Weckesser, 2019). For example, with $\lambda<-14$ and $x>1$, we have $x^{\lambda}\to 0$, and the subtraction $x^{\lambda}-1$ leads to severe precision loss. Therefore, computing the variance after the transformation becomes very unstable. The same occurs for large $\lambda$ with $x<1$. This instability, illustrated in the left panel of Figure 5, shows large fluctuations in the NLL curve when keeping the constant term, which will disrupt optimization whenever the evaluated $\lambda$ lies in such regions.
@@
Here, $\psi^{-1}_{\text{BC}}$ is the inverse Box-Cox transform, expressed via the Lambert $W$ function (Corless et al., 1996):
@@
The same approach applies to the Yeo-Johnson transformation (see Appendix C). The default bound $y_{B}$ can be set to the maximum representable floating-point value to prevent overflow after transformation. Practitioners may select a smaller bound based on domain-specific requirements or constraints on downstream models.
@@ 5 Federated Power Transform
To reduce communication rounds, one-pass variance computation is preferred. However, this approach introduces severe numerical instabilities, which directly affect both NLL evaluation and the optimization of $\lambda$. To illustrate this, we compare two approaches: the naive one-pass method (Equation (5), also used in Marchand et al. (2022)) and our numerically stable aggregation procedure (Algorithm 1). As shown in Figure 6, the naive method produces large fluctuations in the NLL curve, disrupting optimization. In contrast, our stable approach yields smooth curves and enables reliable optimization.
@@
Algorithm 2 Federated NLL Comptuation (Box-Cox)
@@ 6 Empirical Evaluation
Our experiments contain two parts: (1) evaluating the effect of power transforms on downstream tasks, and (2) testing the numerical stability of our methods. Code is available at https://github.com/xuefeng-xu/powertf.
@@ 6.1 Downstream Effectiveness
Figure 8: Effect of varying $\lambda$ on AUC scores (LDA).
We first evaluate the impact of power transforms on downstream classification tasks using three datasets: Adult (Becker and Kohavi, 1996), Bank (Moro et al., 2012), and Credit (Yeh, 2009). Dataset statistics are shown in Table 3. Each dataset is split into 80% training and 20% testing. Since features may contain negative values, we apply the Yeo-Johnson transformation to each feature, estimating $\lambda^{*}$ from the training set. We compare against two baselines: (1) Standardization (STD: zero mean and unit variance), and (2) Raw data without any transformation.
We then train three classifiers on the transformed features: Linear Discriminant Analysis (LDA), Logistic Regression (LR), and XGBoost (XGB). Since LDA assumes normally distributed inputs, power transforms are expected to improve performance. Figure 7 shows ROC curves and AUC scores on the test set (additional plots are in Figure 11 in the Appendix). Power transforms consistently outperform the baselines, although the improvements are modest. Notably, for XGBoost, the performance differential between using raw and preprocessed data is negligible. A recent study by Eftekhari and Papyan (2025) applied power transforms to hidden layers of deep neural networks, showing that improved Gaussianity leads to higher classification accuracy. This further confirms the effectiveness of power transforms.
We next study the effect of varying $\lambda$ on AUC scores. This tests whether the estimated $\lambda^{*}$ is indeed optimal for downstream tasks. Since each dataset has multiple features, we select the feature with the highest mutual information with the label and ignore other features and apply the Yeo-Johnson transform with varying $\lambda$. Figures 8 shows AUC as a function of $\lambda$ (additional results are in Figure 12 in the Appendix).